### v1

# Efficient Log-Rank Updates for Random Survival Forests

Abstract
Random survival forests are widely used for estimating covariate-conditional survival functions under right-censoring. Their standard log-rank splitting criterion is typically recomputed at each candidate split. This $O(M)$ cost per split, with $M$ the number of distinct event times in a node, creates a bottleneck for large cohort datasets with long follow-up. We revisit approximations proposed by LeBlanc and Crowley (1995) and develop simple constant-time updates for the log-rank criterion. The method is implemented in grf and substantially reduces training time on large datasets while preserving predictive performance.
1 Introduction
Random forests (Breiman, 2001) are a staple of the machine learning toolbox, often excelling at conditional mean prediction tasks on tabular data. Since their introduction, Breiman’s algorithmic blueprint has been extended beyond regression to a wide range of statistical quantities. For survival analysis, Ishwaran et al. (2008) introduced random survival forests to estimate the conditional survival function $S(t;x)=\mathbb{P}\left[T_{i}>t\mid X_{i}=x\right]$. A random survival forest typically proceeds in two phases. First, the covariate space is recursively partitioned via axis-aligned cuts, with splits chosen to maximize a criterion that separates units into groups with differing survival. Second, the resulting partitions are used to produce covariate-specific predictions, either by ensemble-aggregating or kernel-weighting Kaplan–Meier (or Nelson–Aalen) survival functions. The most common splitting criterion is the log-rank statistic, initially suggested for survival trees by LeBlanc and Crowley (1993); Segal (1988). Intuitively, the log-rank statistic assesses whether a candidate split separates samples into regions with different survival distributions.
A major reason for the computational success of Breiman’s regression forests is the efficiency of CART splitting, which reduces to computing running means (Hastie et al., 2009). Given sorted samples, evaluating all $n$ candidate split points takes $O(n)$ time. Similar efficiency gains underlie many practical extensions of forests, such as the use of gradient-based approximations in boosted trees (Friedman, 2001) or the generalized random forest framework of Athey et al. (2019), where custom targets defined via estimating equations are reduced to CART updates. In contrast, random survival forests do not typically admit such simplifications: the log-rank statistic requires summing over the $M$ distinct event times in a node, giving $O(M)$ computation for every candidate split. To the best of our knowledge, leading implementations, including randomForestSRC (Ishwaran et al., 2008), ranger (Wright and Ziegler, 2017), and sksurv (Pölsterl, 2020), all implement log-rank splitting in this way [^1] The respective code is here for randomForestSRC, here for ranger, and here for sksurv., as do other approaches relying on two-sample tests for split evaluation (Hothorn and Lausen, 2003).
These computational demands are especially challenging for large datasets with long follow-up, now common in modern cohort studies. For example, the SEER cancer database (Howlader et al., 2021) follows patients for several decades and records numerous high-cardinality predictors (e.g., continuous measures of mitotic rate). In such settings, modern statistical learning methods are increasingly deployed to estimate nuisance components for causal quantities (Wager, 2024). For instance, when estimating heterogeneous treatment effects in time-to-event data with right-censoring, statistical efficiency requires accounting for censoring through doubly robust adjustments (Cui et al., 2023). These adjustments are necessary to obtain unbiased estimation in the presence of censoring, but they require estimating two conditional survival functions: one for the event process and one for the censoring process.
To mitigate computational burden, current practice often relies on heuristics, such as restricting the number of events considered, subsampling split points, or screening covariates. Motivated by this bottleneck, we seek an algorithmic solution closer in spirit to Breiman’s original one-pass scanning step by leveraging approximations. Deriving efficient update rules for the log-rank statistic is complicated by its variance calculation. In early work studying survival trees, LeBlanc and Crowley (1993) noted that such updates are possible for the numerator of the statistic, while LeBlanc and Crowley (1995) later noted variance approximations that could enable constant-time updates of the denominator.
To our knowledge, these observations have not been pursued in subsequent methodology or software, perhaps due to the perceived complexity of variance updates. In this note, we formalize these approximations into simple constant-time update rules that can be readily implemented in existing software. This simple algorithmic refinement is implemented in the survival_forest function of the grf (Tibshirani et al., 2025) package for (R Core Team, 2024). On a simulated example with 100,000 units observed on a 260-month time grid ($\approx$ 20 years), and 50 continuous predictors, an optimized forest runs about $3\times$ faster, while maintaining essentially identical statistical performance.
2 Log-rank splitting
To ease exposition, we restrict ourselves to the salient part of the random forest algorithm where the computation of the log-rank criterion plays an important role. We restrict our attention to a given node containing $n$ samples measuring the outcomes $T_{i}\in\{1,2,\ldots,M\}$ that record some time point $t=1,\ldots,M$ [^2] We can easily obtain the $M$ unique event times in any candidate node and remap them to consequtive integers $1,\ldots$ via binary search.. We also have access to the event indicator variable $D_{i}\in\{0,1\}$ indicating if the $i$-th sample is censored ($D_{i}=0$) or observed ($D_{i}=1$; this is often referred to as a “failure time” or “death”). $M$ is the number of time points for which $D_{i}=1$. For these $n$ samples, we consider a single candidate predictor variable $X$, which we take as given in sorted order $X_{1}<X_{2}\ldots<X_{n}$. For a given candidate split point $X_{i}=c$, let
$L=\{i\mid X_{i}\leq c\}$
denote the set of samples in the left node and
$R=\{i\mid X_{i}>c\}$
the set of samples in the right node. Let $d_{t}$ denote the number of samples with a failure event ($D_{i}=1$) at the $t$-th time point:
$d_{t}=|\{i\mid T_{i}=t,D_{i}=1\}|,$ (1)
where we use the subscripts $L,R$ to denote the left and right node count, with:
$d_{t,L}+d_{t,R}=d_{t}.$
Let $Y_{t}$ denote the number “at risk” at the $t$-th timepoint
$Y_{t}=|\{i\mid T_{i}\geq t\}|,$ (2)
with the same left-right counterparts:
$Y_{t,L}+Y_{t,R}=Y_{t}.$
The log-rank criterion at the candidate split value $c$ is (as implemented and documented in Ishwaran and Kogalur (2025))
$\mathcal{L}(c)=\frac{\sum\limits_{t=1}^{M}(d_{t,L}-Y_{t,L}\alpha_{t})}{\sqrt{\sum\limits_{t=1}^{M}Y_{t,L}(Y_{t}-Y_{t,L})\beta_{t}}},$ (3)
where we have defined the node-specific coefficients $\alpha_{t}$ and $\beta_{t}$ that stay constant for different split points $c$,
$\alpha_{t}=\frac{d_{t}}{Y_{t}},$ (4)
$\beta_{t}=\left(\frac{Y_{t}-d_{t}}{Y_{t}-1}\right)\frac{d_{t}}{Y_{t}^{2}}.$ (5)
Conceptually, this is a simple two-sample test for the null that the expected number of failures in the two nodes is the same, where under the standard sampling model $d_{t,L}$ has mean $Y_{t,L}\alpha_{t}$ and variance $Y_{t,L}(Y_{t}-Y_{t,L})\beta_{t}$.
3 Efficient log-rank splits
A computationally demanding step in a random survival forest is scanning over all candidate split points $c$ to compute $\mathcal{L}(c)^{2}$ and select the split that maximizes it. A direct calculation of (3) requires $O(M)$ operations per candidate, giving this scanning step a total runtime of $O(nM)$ for $n$ samples and $M$ distinct event times. To more closely mimic the linear-time efficiency of Breiman’s original regression forests, we aim to update $\mathcal{L}(c)$ in constant time for each candidate, reducing the overall scan to $O(n)$.
3.1 Exact numerator updates
At first glance, it may appear that the numerator of $\mathcal{L}(c)$, being a sum over $t=1,\dots,M$, must be recomputed in $O(M)$ time for each candidate. However, as noted by LeBlanc and Crowley (1993), the numerator is linear in $d_{t,L}$ and $Y_{t,L}$ and can be updated in $O(1)$ time. This observation enables the linear-time scan over all split points. To see this, consider the first term of the numerator, which, using (1) and changing the order of summation, can be written as
$\sum_{t=1}^{M}d_{t,L}$ $=\sum_{t=1}^{M}\sum_{i=1}^{n}\mathds{1}\!\left\{T_{i}=t\right\}\mathds{1}\!\left\{D_{i}=1\right\}\mathds{1}\!\left\{X_{i}\leq c\right\}$ (6)
$=\sum_{i=1}^{n}\sum_{t=1}^{M}\mathds{1}\!\left\{T_{i}=t\right\}\mathds{1}\!\left\{D_{i}=1\right\}\mathds{1}\!\left\{X_{i}\leq c\right\}$ (7)
$=\sum_{i=1}^{n}D_{i}\mathds{1}\!\left\{X_{i}\leq c\right\}.$ (8)
Similarly, using (2) we get that the second term is
$\sum_{t=1}^{M}Y_{t,L}\alpha_{t}$ $=\sum_{t=1}^{M}\sum_{i=1}^{n}\mathds{1}\!\left\{T_{i}\geq t\right\}\mathds{1}\!\left\{X_{i}\leq c\right\}\alpha_{t}$ (9)
$=\sum_{i=1}^{n}\sum_{t=1}^{M}\mathds{1}\!\left\{T_{i}\geq t\right\}\alpha_{t}\mathds{1}\!\left\{X_{i}\leq c\right\}$ (10)
$=\sum_{i=1}^{n}\gamma_{i}\mathds{1}\!\left\{X_{i}\leq c\right\},$ (11)
where we have defined
$\gamma_{i}=\sum_{t=1}^{M}\mathds{1}\!\left\{T_{i}\geq t\right\}\alpha_{t}.$ (12)
Together, we have that the numerator of (3) at a given split $c$ can be written as
$\mathcal{L}_{\text{num}}(c)=\sum_{i=1}^{n}(D_{i}-\gamma_{i})\mathds{1}\!\left\{X_{i}\leq c\right\}.$
This form depends only on the per-sample quantities $D_{i}$ and $\gamma_{i}$, allowing a single-pass $O(n)$ update over all split points.
3.2 Approximate denominator updates via Poissonization
The denominator of (3) involves hypergeometric variances, which are quadratic in $Y_{t,L}$ and cannot be updated in constant time. A simple approximation (Cox and Oakes, 1984, p. 105) replaces the hypergeometric variance with
$\left(\frac{1}{E_{1}(c)}+\frac{1}{E_{2}(c)}\right)^{-1},$
where $E_{1}(c)$ and $E_{2}(c)$ are the expected number of failures in the left and right nodes. Using the $\gamma_{i}$ from Section 3.1, these expectations can be computed in a single pass:
$E_{1}(c)$ $=\sum_{i=1}^{n}\gamma_{i}\mathds{1}\!\left\{X_{i}\leq c\right\},$ (13)
$E_{2}(c)$ $=\sum_{i=1}^{n}\gamma_{i}\mathds{1}\!\left\{X_{i}>c\right\}=\sum_{i=1}^{n}\gamma_{i}-E_{1}(c)$ (14)
The denominator can then be approximated by
$\mathcal{L}_{\text{den}}(c)\approx\left(\frac{1}{E_{1}(c)}+\frac{1}{E_{2}(c)}\right)^{-\frac{1}{2}},$
which, like the numerator, can be updated in $O(1)$ time per split point.
This approximation is well-motivated. It is well-known that the log-rank statistic is simply the score of a Cox regression. Early work recognized that Cox models could be approximated using a Poisson representation (Laird and Olivier, 1981; Whitehead, 1980), but this insight appears to have been largely overlooked in modern log-rank–based survival forests. LeBlanc and Crowley (1995) exploited this idea by Poissonizing the likelihood to derive fast approximations to the Cox score in a general step-function framework. In special cases, their score reduces to the standard log-rank statistic, with the corresponding Poissonized variance matching the formula suggested by Cox and Oakes (1984) (see LeBlanc and Crowley (1995, Section 3)).
3.3 Efficient survival splitting
Combining the exact numerator update with the approximate denominator gives what we will refer to as the LeBlanc & Crowley splitting criterion for survival forests:
[…]
Algorithm 1 implements these updates, achieving $O(n)$ evaluation of all candidate splits for a single variable333grf has additional splitting logic for missing $X_{i}$-values, which essentially involves running the algorithm twice: once sending all missing values to the left, and once sending all missing values to the right (and finally, evaluating splits on missing, Mayer et al. (2020); Twala et al. (2008)). Practical constraints on minimum node sizes are omitted here..
Input : Samples $X_{i}$, $i=1,\ldots,n$, sorted in increasing order along with times $T_{i}$ and events $D_{i}$ on time index $t=1\ldots M$.
Output : Best approximate log-rank and split value.
Construct $d_{t}$ and $Y_{t}$ vectors
Compute node-specific weights $\alpha_{t}\leftarrow\dfrac{d_{t}}{Y_{t}}$
Initialize cumulative sums $A_{t}\leftarrow\sum\limits_{s=1}^{M}\alpha_{s}\mathds{1}\!\left\{s\leq t\right\}$
Compute node sum $\bar{\gamma}\leftarrow\sum\limits_{t=1}^{M}A_{t}$
Initialize numerator $\mathcal{L}_{\text{num}}\leftarrow 0$
Initialize $E_{1}\leftarrow 0$
[…]
for $i=1\ldots n$ do
   Set $j$ to the $i$-th sample time
   Set $\gamma_{i}\leftarrow A_{j}$
   Update numerator $\mathcal{L}_{\text{num}}\leftarrow\mathcal{L}_{\text{num}}+D_{i}-\gamma_{i}$
   Update variance components
 $E_{1}\leftarrow E_{1}+\gamma_{i}$
 $E_{2}\leftarrow\bar{\gamma}-E_{1}$
   Compute $i$-th squared log-rank statistic with approximate variance
[…]
end for
[…]
To highlight the computational difference between the exact and accelerated splitting rules, we benchmark the time to grow a single survival tree. Since forest training is just an ensemble of such trees, this isolates the cost of evaluating splits and provides a direct comparison of the $O(nM)$ vs. $O(n)$ regimes. For reference, the total forest training time scales linearly with the number of trees. We draw survival times from a Poisson distribution with approximately $M=\{20,130,260,500\}$ distinct events to mimic settings with monthly follow-up ranging from a few months to several decades (including one final setting with a nearly continuous grid, e.g., events measured in days). We set around 10% of the units to be censored. We then vary $p$ continuous covariates with a sample size of $n$, and measure the time it takes to grow a single tree on the full set of samples. Table 1 reports the mean runtimes, showing that $O(1)$ splitting yields a consistent speedup.
$n$ $p$ $M$ Runtime: Exact (s) Runtime: Approx (s) Speedup ($\times$)
20 000 25 20 0.15 0.12 1.26
50 000 25 20 0.46 0.36 1.26
250 000 25 20 3.34 2.77 1.21
20 000 50 20 0.32 0.25 1.28
50 000 50 20 1.00 0.81 1.24
250 000 50 20 6.97 5.70 1.22
20 000 100 20 0.67 0.52 1.28
50 000 100 20 2.08 1.71 1.21
250 000 100 20 14.20 11.62 1.22
20 000 25 130 0.34 0.14 2.50
50 000 25 130 0.94 0.42 2.26
250 000 25 130 5.89 3.05 1.93
20 000 50 130 0.68 0.28 2.46
50 000 50 130 1.96 0.88 2.23
250 000 50 130 11.90 6.18 1.93
20 000 100 130 1.44 0.59 2.44
50 000 100 130 4.05 1.87 2.17
250 000 100 130 23.89 12.35 1.93
20 000 25 260 0.58 0.15 3.90
50 000 25 260 1.56 0.45 3.48
250 000 25 260 8.22 3.20 2.57
20 000 50 260 1.20 0.30 3.96
50 000 50 260 3.23 0.97 3.32
250 000 50 260 17.40 6.50 2.68
20 000 100 260 2.33 0.63 3.68
50 000 100 260 6.45 1.95 3.32
250 000 100 260 35.33 13.24 2.67
20 000 25 500 1.06 0.16 6.49
50 000 25 500 2.75 0.48 5.74
250 000 25 500 13.68 3.44 3.97
20 000 50 500 2.22 0.33 6.64
50 000 50 500 5.58 1.02 5.47
250 000 50 500 27.59 6.98 3.95
20 000 100 500 4.32 0.66 6.51
50 000 100 500 11.82 2.13 5.55
250 000 100 500 57.38 14.05 4.08
Table 1: Run times (in seconds) for growing a single tree (on all $n$ samples) using either the exact log-rank criterion (3) or the approximate log-rank criterion (15) computed using Algorithm 1, across increasing sample size $n$, covariate dimension $p$, and total number of events $M$. The run times are averaged over 10 repetitions.
4 Empirical comparison of approximate and exact splits
Algorithm 1 offers a clear computational advantage on large datasets: updating in $O(1)$ time rather than $O(M)$ can substantially reduce runtime. Equally important, however, is whether this speedup comes without loss of statistical performance.
4.1 Comparing concordance on benchmark data sets
Following Ishwaran et al. (2008), we measure prediction error $PE_{C}$ as the complement of the CC–index (Harrell et al., 1982) with the conditional cumulative hazard as outcomes (computed from the estimated Nelson–Aalen curves). This metric ranges from 0 (perfect concordance) to 1 (worst), with 0.5 corresponding to random chance. We benchmark performance on survival data sets from the survival package (Therneau, 2024): Stanford heart transplant (heart); NCCTG lung cancer (lung); Mayo Clinic primary biliary cholangitis (pbc); breast cancer from the Rotterdam tumor bank (rotterdam); and the Veterans Administration lung cancer study (veteran). For each dataset, we compute the paired difference in prediction error, $\Delta PE_{C}=PE_{C}(\text{exact})-PE_{C}(\text{approx})$, by fitting random survival forests under identical settings and seeds, once with exact log-rank splits and once with approximate splits. Figure 1 shows boxplots of $\Delta PE_{C}$ over 250 random seeds. Across all data sets, $\Delta PE_{C}$ is negligible, between $10^{-4}$ and $5\times 10^{-3}$ (an order of magnitude smaller than what’s usually considered meaningful in survival model comparisons), indicating that the two algorithms are effectively indistinguishable in predictive accuracy.
Figure 1: Boxplots of difference in prediction error $\Delta PE_{C}=PE_{C}(\text{exact})-PE_{C}(\text{approx})$ over 250 repetitions. Typically, good survival models achieve concordance in the 0.250.350.25-0.35 range. The prediction error difference here is orders of magnitude smaller than this baseline.
4.2 Comparing RMSE in simulated settings
We adapt the four simulation settings considered in Cui et al. (2023) (and add a fifth) to estimate a given time point $h$ on the conditional survival function $S(t;x)$. We measure prediction error $PE_{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(S(h;X_{i})-\widehat{S}(h;X_{i})\right)^{2}}$ with out-of-bag estimates. The number of samples considered is $n=5000$ with $p=10$ covariates. For details on the data generating processes, including the query time point $h$ for each setup, we refer to Cui et al. (2023, Section 4). For each simulation setting, we compute the paired difference in prediction error, $\Delta PE_{RMSE}=PE_{RMSE}(\text{exact})-PE_{RMSE}(\text{approx})$, as in Section 4.1. Figure 2 shows boxplots of $\Delta PE_{RMSE}$ over 250 simulation repetitions. Across all five simulation settings, the difference in RMSE between the exact and approximate criteria is on the order of $10^{-3}$, which is negligible compared to the typical RMSE values (typically measured in several percent for survival probabilities).
Figure 2: Boxplots of difference in prediction error $\Delta PE_{RMSE}=PE_{RMSE}(\text{exact})-PE_{RMSE}(\text{approx})$ over 250 repetitions. The difference in RMSE indicates an average change in predicted survival probabilities on the order of roughly one-hundredth of a percent, effectively negligible.
5 Discussion
Early survival tree methods were motivated by two-sample tests such as the log-rank, designed with power considerations for a single hypothesis test in a clinical setting with right-censoring. In the context of modern machine learning, and especially Breiman’s random forests, the goal is different: identify directions in covariate space that effectively separate survival curves, rather than to optimize a single test. Making the algorithm computationally feasible necessarily involves approximations (e.g., Athey et al. (2019); Friedman (2001)), and because splits are chosen repeatedly across many trees and subsamples, the ensemble averages out local imperfections, so an approximate criterion performs essentially as well as the exact one.
Acknowledgments
We thank Ayush Kanodia for many enlightening algorithm discussions. We also thank Max Schuessler for valuable insights regarding long-term survival data in observational cancer studies.