### v2

# Efficient Log-Rank Updates for Random Survival Forests

Abstract
Random survival forests are widely used for estimating covariate-conditional survival functions under right-censoring. Their standard log-rank splitting criterion is typically recomputed at each candidate split. This $O(M)$ cost per split, with $M$ the number of distinct event times in a node, creates a bottleneck for large cohort datasets with long follow-up. We revisit approximations proposed by LeBlanc and Crowley (1995) and develop simple constant-time updates for the log-rank criterion. The method is implemented in grf for R and reduces training time on large datasets while preserving predictive accuracy.
1 Introduction
Random forests (Breiman, 2001) are a staple of the machine learning toolbox, often excelling at conditional mean prediction tasks on tabular data. Since their introduction, Breiman’s algorithmic blueprint has been extended beyond regression to a wide range of statistical quantities. For survival analysis, Ishwaran et al. (2008) introduced random survival forests to estimate the conditional survival function $S(t;x)=\mathbb{P}\left[T_{i}>t\mid X_{i}=x\right]$. A random survival forest typically proceeds in two phases. First, the covariate space is recursively partitioned via axis-aligned cuts, with splits chosen to maximize a criterion that separates units into groups with differing survival. Second, the resulting partitions are used to produce covariate-specific predictions, either by ensemble-aggregating or kernel-weighting Kaplan–Meier (or Nelson–Aalen) survival functions. The most common splitting criterion is the log-rank statistic, initially suggested for survival trees by LeBlanc and Crowley (1993); Segal (1988). Intuitively, the log-rank statistic assesses whether a candidate split separates samples into regions with different survival distributions.
A major reason for the computational success of Breiman’s regression forests is the efficiency of CART splitting, which reduces to computing running means (Hastie et al., 2009). Given sorted samples, evaluating all $n$ candidate split points takes $O(n)$ time. Similar efficiency gains underlie many practical extensions of forests, such as the use of gradient-based approximations in boosted trees (Friedman, 2001) or the generalized random forest framework of Athey et al. (2019), where custom targets defined via estimating equations are reduced to CART splits on gradient-based pseudo outcomes. In contrast, random survival forests do not typically admit such simplifications: the log-rank statistic requires summing over the $M$ distinct event times in a node, giving $O(M)$ computation for every candidate split. To the best of our knowledge, leading implementations, including randomForestSRC (Ishwaran et al., 2008), ranger (Wright and Ziegler, 2017), and sksurv (Pölsterl, 2020), all implement log-rank splitting in this way, as do other approaches relying on two-sample tests for split evaluation (Hothorn and Lausen, 2003).
These computational demands are especially challenging for large datasets with long follow-up, common in modern cohort studies. For example, the SEER cancer database (Howlader et al., 2021) follows patients for several decades and records numerous high-cardinality predictors, such as continuous measures of mitotic rate. In these settings, modern statistical learning methods are increasingly deployed to estimate nuisance components for causal quantities (see, e.g., Wager, 2024). For instance, when estimating heterogeneous treatment effects or regression discontinuity parameters in time-to-event data with right-censoring (e.g., Cui et al. (2023); Schuessler et al. (2026)), accounting for censoring is necessary for unbiased estimation. When flexible machine learning estimators, such as random survival forests, are used, this is typically achieved via doubly robust censoring adjustments (Robins et al., 1994; Rubin and van der Laan, 2007). These adjustments require estimating two conditional survival functions: one for the event process and one for the censoring process.
To mitigate computational burden, current practice often relies on heuristics, such as restricting the number of events considered, subsampling split points, or screening covariates. Motivated by this bottleneck, we seek an algorithmic solution closer in spirit to Breiman’s original one-pass scanning step by leveraging approximations. Deriving efficient update rules for the log-rank statistic is complicated by its variance calculation. In early work studying survival trees, LeBlanc and Crowley (1993) noted that such updates are possible for the numerator of the statistic, while LeBlanc and Crowley (1995) later noted variance approximations that could enable constant-time updates of the denominator.
To our knowledge, these observations have not been pursued in subsequent methodology or software, perhaps due to the perceived complexity of variance updates. In this note, we formalize these approximations into simple constant-time update rules that can be readily implemented in existing software. This simple algorithmic refinement is implemented in the survival_forest function of the grf (Tibshirani et al., 2025) package for (R Core Team, 2024), and controlled with the fast.logrank argument. On a simulated example with 100 000 units observed on a 260-month time grid ($\approx$ 20 years), and 50 continuous predictors, an optimized forest runs about $3\times$ faster, while maintaining essentially identical statistical performance.
2 Log-rank splitting
To ease exposition, we restrict ourselves to the salient part of the random forest algorithm where the computation of the log-rank criterion plays an important role. We restrict our attention to a given node containing $n$ samples measuring the outcomes $T_{i}\in\{1,2,\ldots,M\}$ that record some time point $t=1,\ldots,M$ [^1] We can easily obtain the $M$ unique event times in any candidate node and remap them to consecutive integers $1,\ldots$ via binary search.. We also have access to the event indicator variable $D_{i}\in\{0,1\}$ indicating if the $i$-th sample is censored ($D_{i}=0$) or observed ($D_{i}=1$; this is often referred to as a “failure time” or “death”). $M$ is the number of time points for which $D_{i}=1$. For these $n$ samples, we consider a single candidate predictor variable $X$, which we take as given in sorted order $X_{1}<X_{2}\cdots<X_{n}$. For a given candidate split point $X_{i}=c$, let
$L=\{i\mid X_{i}\leq c\}$
denote the set of samples in the left node and
$R=\{i\mid X_{i}>c\}$
the set of samples in the right node. Let $d_{t}$ denote the number of samples with a failure event ($D_{i}=1$) at the $t$-th time point:
$d_{t}=|\{i\mid T_{i}=t,D_{i}=1\}|,$ (1)
where we use the subscripts $L,R$ to denote the left and right node count, with:
$d_{t,L}+d_{t,R}=d_{t}.$
Let $Y_{t}$ denote the number “at risk” at the $t$-th timepoint
$Y_{t}=|\{i\mid T_{i}\geq t\}|,$ (2)
with the same left-right counterparts:
$Y_{t,L}+Y_{t,R}=Y_{t}.$
The log-rank criterion at the candidate split value $c$ is (as implemented and documented in Ishwaran and Kogalur (2025))
$\mathcal{L}(c)=\frac{\sum\limits_{t=1}^{M}(d_{t,L}-Y_{t,L}\alpha_{t})}{\sqrt{\sum\limits_{t=1}^{M}Y_{t,L}(Y_{t}-Y_{t,L})\beta_{t}}},$ (3)
where we have defined the node-specific coefficients $\alpha_{t}$ and $\beta_{t}$ that stay constant for different split points $c$,
$\alpha_{t}=\frac{d_{t}}{Y_{t}},$ (4)
$\beta_{t}=\left(\frac{Y_{t}-d_{t}}{Y_{t}-1}\right)\frac{d_{t}}{Y_{t}^{2}}.$ (5)
Conceptually, this is a simple two-sample test for the null that the expected number of failures in the two nodes is the same, where under the standard sampling model $d_{t,L}$ has mean $Y_{t,L}\alpha_{t}$ and variance $Y_{t,L}(Y_{t}-Y_{t,L})\beta_{t}$.
3 Efficient log-rank splits
A computationally demanding step in a random survival forest is scanning over all candidate split points $c$ to compute $\mathcal{L}(c)^{2}$ and select the split that maximizes it. A direct calculation of (3) requires $O(M)$ operations per candidate, giving this scanning step a total runtime of $O(nM)$ for $n$ samples and $M$ distinct event times. To more closely mimic the linear-time efficiency of Breiman’s original regression forests, we aim to update $\mathcal{L}(c)$ in constant time for each candidate, reducing the overall scan to $O(n)$.
3.1 Exact numerator updates
At first glance, it may appear that the numerator of $\mathcal{L}(c)$, being a sum over $t=1,\dots,M$, must be recomputed in $O(M)$ time for each candidate. However, as noted by LeBlanc and Crowley (1993), the numerator is linear in $d_{t,L}$ and $Y_{t,L}$ and can be updated in $O(1)$ time. This observation enables the linear-time scan over all split points. To see this, consider the first term of the numerator, which, using (1) and changing the order of summation, can be written as
$\sum_{t=1}^{M}d_{t,L}$ $=\sum_{t=1}^{M}\sum_{i=1}^{n}\mathds{1}\!\left\{T_{i}=t\right\}\mathds{1}\!\left\{D_{i}=1\right\}\mathds{1}\!\left\{X_{i}\leq c\right\}$ (6)
$=\sum_{i=1}^{n}\sum_{t=1}^{M}\mathds{1}\!\left\{T_{i}=t\right\}\mathds{1}\!\left\{D_{i}=1\right\}\mathds{1}\!\left\{X_{i}\leq c\right\}$ (7)
$=\sum_{i=1}^{n}D_{i}\mathds{1}\!\left\{X_{i}\leq c\right\}.$ (8)
Similarly, using (2) we get that the second term is
$\sum_{t=1}^{M}Y_{t,L}\alpha_{t}$ $=\sum_{t=1}^{M}\sum_{i=1}^{n}\mathds{1}\!\left\{T_{i}\geq t\right\}\mathds{1}\!\left\{X_{i}\leq c\right\}\alpha_{t}$ (9)
$=\sum_{i=1}^{n}\sum_{t=1}^{M}\mathds{1}\!\left\{T_{i}\geq t\right\}\alpha_{t}\mathds{1}\!\left\{X_{i}\leq c\right\}$ (10)
$=\sum_{i=1}^{n}\gamma_{i}\mathds{1}\!\left\{X_{i}\leq c\right\},$ (11)
where we have defined
$\gamma_{i}=\sum_{t=1}^{M}\mathds{1}\!\left\{T_{i}\geq t\right\}\alpha_{t}.$ (12)
Together, we have that the numerator of (3) at a given split $c$ can be written as
$\mathcal{L}_{\text{num}}(c)=\sum_{i=1}^{n}(D_{i}-\gamma_{i})\mathds{1}\!\left\{X_{i}\leq c\right\}.$
This form depends only on the per-sample quantities $D_{i}$ and $\gamma_{i}$, allowing a single-pass $O(n)$ update over all split points.
3.2 Approximate denominator updates via Poissonization
The denominator of (3) involves hypergeometric variances, which are quadratic in $Y_{t,L}$ and cannot be updated in constant time. A simple approximation (Cox and Oakes, 1984, p. 105) replaces the hypergeometric variance with
$\left(\frac{1}{E_{1}(c)}+\frac{1}{E_{2}(c)}\right)^{-1},$
where $E_{1}(c)$ and $E_{2}(c)$ are the expected number of failures in the left and right nodes. Using the $\gamma_{i}$ from Section 3.1, these expectations can be computed in a single pass:
$E_{1}(c)$ $=\sum_{i=1}^{n}\gamma_{i}\mathds{1}\!\left\{X_{i}\leq c\right\},$ (13)
$E_{2}(c)$ $=\sum_{i=1}^{n}\gamma_{i}\mathds{1}\!\left\{X_{i}>c\right\}=\sum_{i=1}^{n}\gamma_{i}-E_{1}(c)$ (14)
The denominator can then be approximated by
$\mathcal{L}_{\text{den}}(c)\approx\left(\frac{1}{E_{1}(c)}+\frac{1}{E_{2}(c)}\right)^{-\frac{1}{2}},$
which, like the numerator, can be updated in $O(1)$ time per split point.
This approximation is well-motivated. It is well-known that the log-rank statistic is simply the score of a Cox regression. Early work recognized that Cox models could be approximated using a Poisson representation (Laird and Olivier, 1981; Whitehead, 1980), but this insight appears to have been largely overlooked in modern log-rank–based survival forests. LeBlanc and Crowley (1995) exploited this idea by Poissonizing the likelihood to derive fast approximations to the Cox score in a general step-function framework. In special cases, their score reduces to the standard log-rank statistic, with the corresponding Poissonized variance matching the formula suggested by Cox and Oakes (1984) (see LeBlanc and Crowley (1995, Section 3)).
3.3 Efficient survival splitting
Combining the exact numerator update with the approximate denominator gives what we will refer to as the LeBlanc & Crowley splitting criterion for survival forests:
[…]
Algorithm 1 implements these updates, achieving $O(n)$ evaluation of all candidate splits for a single variable222grf has additional splitting logic for missing $X_{i}$-values, which essentially involves running the algorithm twice: once sending all missing values to the left, and once sending all missing values to the right (and finally, evaluating splits on missing, Mayer et al. (2020); Twala et al. (2008)). Practical constraints on minimum node sizes are omitted here..
Input : Samples $X_{i}$, $i=1,\ldots,n$, sorted in increasing order along with times $T_{i}$ and events $D_{i}$ on time index $t=1\ldots M$.
Output : Best approximate log-rank and split value.
Construct $d_{t}$ and $Y_{t}$ vectors
Compute node-specific weights $\alpha_{t}\leftarrow\dfrac{d_{t}}{Y_{t}}$
Initialize cumulative sums $A_{t}\leftarrow\sum\limits_{s=1}^{M}\alpha_{s}\mathds{1}\!\left\{s\leq t\right\}$
Compute node sum $\bar{\gamma}\leftarrow\sum\limits_{t=1}^{M}A_{t}$
Initialize numerator $\mathcal{L}_{\text{num}}\leftarrow 0$
Initialize $E_{1}\leftarrow 0$
for $i=1\ldots n$ do
   Set $j$ to the $i$-th sample time
   Set $\gamma_{i}\leftarrow A_{j}$
   Update numerator $\mathcal{L}_{\text{num}}\leftarrow\mathcal{L}_{\text{num}}+D_{i}-\gamma_{i}$
   Update variance components
 $E_{1}\leftarrow E_{1}+\gamma_{i}$
 $E_{2}\leftarrow\bar{\gamma}-E_{1}$
   Compute $i$-th squared log-rank statistic with approximate variance
[…]
end for
[…]
To highlight the computational difference between the exact and accelerated splitting rules, we benchmark the time to grow a single survival tree. Since forest training is just an ensemble of such trees, this isolates the cost of evaluating splits and provides a direct comparison of the $O(nM)$ vs. $O(n)$ regimes. For reference, the total forest training time scales linearly with the number of trees. We draw survival times from a Poisson distribution with approximately $M=\{20,130,260,500\}$ distinct events to mimic settings with monthly follow-up ranging from a few months to several decades (including one final setting with a nearly continuous grid, e.g., events measured in days). We set around 10% of the units to be censored. We then vary $p$ continuous covariates with a sample size of $n$, and measure the time it takes to grow a single tree on the full set of samples. Table 1 reports the mean runtimes, showing that $O(1)$ splitting yields a consistent speedup [^3] Fitting deeper trees by relaxing the constraint on the number of events in each child node in grf can lead to slightly larger relative speedups and more noticeable absolute speedups; see Table 3 in the appendix..
As noted in the introduction, a common heuristic is to restrict the number of event times by discretizing the time grid to a small value (e.g., $M=20$). Table 1 shows that this is effective, as increasing $M$ substantially raises training time for the exact splitting rule at fixed $n$ and $p$, albeit at the cost of reduced temporal resolution. In contrast, the approximate criterion is essentially insensitive to $M$, allowing one to retain the full time resolution with little additional computational cost.
$n$ $p$ $M$ Runtime: Exact (s) Runtime: Approx (s) Speedup ($\times$)
20 000 25 20 0.15 0.12 1.26
50 000 25 20 0.46 0.36 1.26
250 000 25 20 3.34 2.77 1.21
20 000 50 20 0.32 0.25 1.28
50 000 50 20 1.00 0.81 1.24
250 000 50 20 6.97 5.70 1.22
20 000 100 20 0.67 0.52 1.28
50 000 100 20 2.08 1.71 1.21
250 000 100 20 14.20 11.62 1.22
20 000 25 130 0.34 0.14 2.50
50 000 25 130 0.94 0.42 2.26
250 000 25 130 5.89 3.05 1.93
20 000 50 130 0.68 0.28 2.46
50 000 50 130 1.96 0.88 2.23
250 000 50 130 11.90 6.18 1.93
20 000 100 130 1.44 0.59 2.44
50 000 100 130 4.05 1.87 2.17
250 000 100 130 23.89 12.35 1.93
20 000 25 260 0.58 0.15 3.90
50 000 25 260 1.56 0.45 3.48
250 000 25 260 8.22 3.20 2.57
20 000 50 260 1.20 0.30 3.96
50 000 50 260 3.23 0.97 3.32
250 000 50 260 17.40 6.50 2.68
20 000 100 260 2.33 0.63 3.68
50 000 100 260 6.45 1.95 3.32
250 000 100 260 35.33 13.24 2.67
20 000 25 500 1.06 0.16 6.49
50 000 25 500 2.75 0.48 5.74
250 000 25 500 13.68 3.44 3.97
20 000 50 500 2.22 0.33 6.64
50 000 50 500 5.58 1.02 5.47
250 000 50 500 27.59 6.98 3.95
20 000 100 500 4.32 0.66 6.51
50 000 100 500 11.82 2.13 5.55
250 000 100 500 57.38 14.05 4.08
Table 1: Runtimes (in seconds) for growing a single tree (on all $n$ samples) using either the exact log-rank criterion (3) or the approximate log-rank criterion (15) computed using Algorithm 1, across increasing sample size $n$, covariate dimension $p$, and total number of events $M$. The runtimes are averaged over 10 repetitions.
4 Empirical comparison of approximate and exact splits
Algorithm 1 offers a clear computational advantage on large datasets: updating in $O(1)$ time rather than $O(M)$ can substantially reduce runtime. Equally important, however, is whether this speedup comes without loss of statistical performance.
4.1 Comparing performance on benchmark data sets
We consider prediction metrics that assess both discrimination and calibration. To compare discrimination, following Ishwaran et al. (2008), we measure prediction error $PE_{C}$ as the complement of the CC–index (Harrell et al., 1982) with the conditional cumulative hazard as outcomes (computed from the estimated Nelson–Aalen curves). This metric ranges from 0 (perfect concordance) to 1 (worst), with 0.5 corresponding to random chance. To compare calibration, we measure prediction error with the integrated Brier score $PE_{IBS}$ (Graf et al., 1999). This metric also ranges from 0 to 1, with 0 denoting perfect calibration.
We benchmark performance on standard survival data sets from the survival package (Therneau, 2024): Stanford heart transplant (heart); NCCTG lung cancer (lung); Mayo Clinic primary biliary cholangitis (pbc); breast cancer from the Rotterdam tumor bank (rotterdam); and the Veterans Administration lung cancer study (veteran). For each dataset, we compute the paired difference in prediction errors, $\Delta PE_{C}=PE_{C}(\text{exact})-PE_{C}(\text{approx})$ and $\Delta PE_{IBS}=PE_{IBS}(\text{exact})-PE_{IBS}(\text{approx})$, by fitting random survival forests under identical settings and seeds, once with exact log-rank splits and once with approximate splits. Figure 1 shows boxplots of $\Delta PE_{C}$ and $\Delta PE_{IBS}$ over 250 random seeds. Across all data sets, $\Delta PE_{C}$ is negligible, between $10^{-4}$ and $5\times 10^{-3}$ (an order of magnitude smaller than what’s usually considered meaningful in survival model comparisons), and similarly for $\Delta PE_{IBS}$, indicating that the two algorithms are effectively indistinguishable in predictive accuracy. Table 4 in the Appendix summarizes results on a larger collection of benchmark datasets from SurvSet (Drysdale, 2022). Across datasets and metrics, the average error difference is only around 0.00005, consistent with the results above.
Figure 1: Boxplots of difference in 1(a) concordance $\Delta PE_{C}=PE_{C}(\text{exact})-PE_{C}(\text{approx})$ and 1(b) integrated Brier score $\Delta PE_{IBS}=PE_{IBS}(\text{exact})-PE_{IBS}(\text{approx})$, over 250 repetitions. Typically, good survival models achieve concordance in the 0.250.350.25-0.35 range and integrated Brier score in the range 0.10.30.1-0.3 The prediction error difference here is orders of magnitude smaller than this baseline.
4.2 Comparing RMSE in simulated settings
We adapt the four simulation settings considered in Cui et al. (2023) (and add a fifth) to estimate a given time point $h$ on the conditional survival function $S(t;x)$. We measure prediction error $PE_{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(S(h;X_{i})-\widehat{S}(h;X_{i})\right)^{2}}$ with out-of-bag estimates. The number of samples considered is nn=5 000 with $p=10$ covariates. For details on the data generating processes, including the query time point $h$ for each setup, we refer to Cui et al. (2023, Section 4). For each simulation setting, we compute the paired difference in prediction error, $\Delta PE_{RMSE}=PE_{RMSE}(\text{exact})-PE_{RMSE}(\text{approx})$, as in Section 4.1. Figure 2 shows boxplots of $\Delta PE_{RMSE}$ over 250 simulation repetitions. Across all five simulation settings, the difference in RMSE between the exact and approximate criteria is on the order of $10^{-3}$, which is negligible compared to the typical RMSE values (typically measured in several percent for survival probabilities).
Figure 2: Boxplots of difference in prediction error $\Delta PE_{RMSE}=PE_{RMSE}(\text{exact})-PE_{RMSE}(\text{approx})$ over 250 repetitions. The difference in RMSE indicates an average change in predicted survival probabilities on the order of roughly one-hundredth of a percent, effectively negligible.
For completeness, we also compare runtimes on these smaller datasets in Table 2. The datasets in survival are small (except for rotterdam, which has over 2 000 distinct event times for roughly 3 000 units) and timing differences are negligible. For the simulated datasets we see similar relative speedups as in Table 1, though at this scale the absolute time differences are unlikely to be noticeable.
Dataset/simulation Runtime: Exact (ms) Runtime: Approx (ms) Speedup ($\times$)
heart 1.40 1.40 1.00
lung 1.40 1.40 1.00
pbc 1.40 1.40 1.00
rotterdam 8.57 3.98 2.15
veteran 1.40 1.38 1.01
Setting 1 31.87 10.47 3.04
Setting 2 31.39 9.47 3.31
Setting 3 13.24 10.08 1.31
Setting 4 9.98 8.71 1.15
Setting 5 31.33 10.00 3.13
Table 2: Runtimes (in milliseconds, averaged over 100 repetitions) for growing a single tree on the datasets used in Figure 1 and Figure 2.
5 Discussion
Early survival tree methods were motivated by two-sample tests such as the log-rank statistic, designed for easily described prognostic subgroups in clinical settings with right-censored outcomes. In contrast, modern random forests aim to identify partitions of the covariate space that effectively separate survival distributions, rather than to optimize a single test statistic.
Approximate criteria based on first-order information have proven effective in related settings, such as gradient boosting (Friedman, 2001) and generalized random forests (Athey et al., 2019). Although the log-rank statistic is not explicitly gradient-based, its numerator can be viewed as a score-like quantity closely related to martingale residuals. Consistent with this, we find that splitting on the numerator $\mathcal{L}_{\text{num}}^{2}$ alone yields similar empirical performance. For the extended benchmarks in Appendix Table 4, the overall mean difference in prediction error between the exact log-rank criterion and this simplified version increases more than tenfold, but remains small in absolute magnitude. Further improvement is obtained by standardizing $\mathcal{L}_{\text{num}}^{2}$ by the product of the fractions of units sent left and right, $n_{L}(c)n_{R}(c)$. This scaling more closely approximates the Poisson variance formulation, which is equivalent to replacing these fractions with the expected number of events in each node.
Prior work has also noted connections between log-rank splitting and least-squares splitting based on martingale residuals (Keleş and Segal, 2002; LeBlanc and Crowley, 1992), further suggesting a connection to gradient-based criteria. From a practical perspective, we prefer the approximate variance–scaled log-rank formulation in grf, as empirical differences are negligible and preserve compatibility with prior work using the exact criterion.
While a formal analysis is beyond the scope of this note, it would be interesting to characterize when the exact and approximate log-rank statistics induce the same rankings, and to what extent the numerator drives the split discrimination.
Acknowledgments
We thank Ayush Kanodia for many enlightening algorithm discussions and Max Schuessler for valuable insights regarding long-term survival data in observational cancer studies. We are also grateful to two anonymous referees for helpful feedback.
Appendix A Additional results
$n$ $p$ $M$ Runtime: Exact (s) Runtime: Approx (s) Speedup ($\times$)
20 000 25 20 0.74 0.51 1.46
50 000 25 20 2.23 1.62 1.38
250 000 25 20 16.02 11.73 1.37
20 000 50 20 1.75 1.34 1.31
50 000 50 20 5.70 4.20 1.36
250 000 50 20 37.89 28.95 1.31
20 000 100 20 4.73 3.71 1.27
50 000 100 20 14.90 11.32 1.32
250 000 100 20 102.38 81.87 1.25
20 000 25 130 1.58 0.55 2.88
50 000 25 130 4.62 1.68 2.75
250 000 25 130 29.41 12.13 2.42
20 000 50 130 3.77 1.36 2.77
50 000 50 130 12.02 4.41 2.73
250 000 50 130 78.35 31.84 2.46
20 000 100 130 10.61 3.77 2.81
50 000 100 130 33.06 12.47 2.65
250 000 100 130 213.85 86.45 2.47
20 000 25 260 2.27 0.57 3.98
50 000 25 260 6.73 1.66 4.06
250 000 25 260 41.18 12.52 3.29
20 000 50 260 5.62 1.37 4.10
50 000 50 260 16.21 4.44 3.65
250 000 50 260 101.84 31.60 3.22
20 000 100 260 14.99 3.84 3.91
50 000 100 260 44.45 12.34 3.60
250 000 100 260 282.84 88.29 3.20
20 000 25 500 3.67 0.60 6.14
50 000 25 500 11.00 1.84 5.98
250 000 25 500 58.12 12.79 4.54
20 000 50 500 9.17 1.41 6.52
50 000 50 500 26.04 4.47 5.83
250 000 50 500 144.47 31.97 4.52
20 000 100 500 22.85 3.93 5.81
50 000 100 500 65.60 12.73 5.15
250 000 100 500 399.65 88.34 4.52
Table 3: Timing results with deeper trees grown under less restrictive constraints on the number of events in each child node (grf option: alpha=0). Runtimes (in seconds) for growing a single tree (on all $n$ samples) using either the exact log-rank criterion or the approximate log-rank criterion, across increasing sample size $n$, covariate dimension $p$, and total number of events $M$. The runtimes are averaged over 5 repetitions.
Split constraint
Dataset n p M Metric Default Relaxed
hdfail 52 422 37 2 641 $\Delta PE_{C}$ 0.00000 0.00004
prostateSurvival 14 294 7 105 $\Delta PE_{C}$ 0.00000 0.00076
support2 9 105 58 1 041 $\Delta PE_{C}$ -0.00036 -0.00066
flchain 7 874 26 1 738 $\Delta PE_{C}$ -0.00002 0.00001
Dialysis 6 805 8 42 $\Delta PE_{C}$ 0.00076 0.00181
dataDIVAT1 5 943 6 810 $\Delta PE_{C}$ 0.00091 0.00117
rhc 5 735 74 731 $\Delta PE_{C}$ 0.00039 0.00042
Framingham 4 699 7 1 372 $\Delta PE_{C}$ -0.00051 0.00030
dataDIVAT3 4 267 7 219 $\Delta PE_{C}$ 0.00000 0.00078
nwtco 4 028 8 392 $\Delta PE_{C}$ 0.00026 -0.00018
smarto 3 873 35 407 $\Delta PE_{C}$ -0.00093 0.00173
acath 3 504 3 244 $\Delta PE_{C}$ 0.00015 -0.00014
divorce 3 371 5 887 $\Delta PE_{C}$ 0.00002 0.00001
UnempDur 3 241 7 26 $\Delta PE_{C}$ -0.00037 -0.00105
rott2 2 982 13 1 078 $\Delta PE_{C}$ -0.00006 -0.00006
Aids2 2 839 13 782 $\Delta PE_{C}$ -0.00049 -0.00081
scania 1 931 6 1 042 $\Delta PE_{C}$ -0.00038 -0.00022
TRACE 1 878 6 958 $\Delta PE_{C}$ -0.00018 0.00011
dataDIVAT2 1 837 4 508 $\Delta PE_{C}$ 0.00026 0.00009
actg 1 151 11 76 $\Delta PE_{C}$ -0.00064 0.00074
LeukSurv 1 043 7 441 $\Delta PE_{C}$ 0.00106 -0.00067
rdata 1 040 7 513 $\Delta PE_{C}$ -0.00037 -0.00025
grace 1 000 5 110 $\Delta PE_{C}$ -0.00006 -0.00031
hdfail 52 422 37 2 641 $\Delta PE_{IBS}$ 0.00000 -0.00046
prostateSurvival 14 294 7 105 $\Delta PE_{IBS}$ 0.00000 0.00002
support2 9 105 58 1 041 $\Delta PE_{IBS}$ -0.00034 -0.00011
flchain 7 874 26 1 738 $\Delta PE_{IBS}$ -0.00001 0.00003
Dialysis 6 805 8 42 $\Delta PE_{IBS}$ 0.00003 0.00025
dataDIVAT1 5 943 6 810 $\Delta PE_{IBS}$ -0.00110 -0.00133
rhc 5 735 74 731 $\Delta PE_{IBS}$ 0.00025 0.00003
Framingham 4 699 7 1 372 $\Delta PE_{IBS}$ -0.00013 0.00007
dataDIVAT3 4 267 7 219 $\Delta PE_{IBS}$ 0.00000 0.00009
nwtco 4 028 8 392 $\Delta PE_{IBS}$ 0.00000 -0.00001
smarto 3 873 35 407 $\Delta PE_{IBS}$ 0.00008 0.00004
acath 3 504 3 244 $\Delta PE_{IBS}$ -0.00009 0.00001
divorce 3 371 5 887 $\Delta PE_{IBS}$ 0.00001 -0.00002
UnempDur 3 241 7 26 $\Delta PE_{IBS}$ -0.00006 -0.00009
rott2 2 982 13 1 078 $\Delta PE_{IBS}$ 0.00040 -0.00020
Aids2 2 839 13 782 $\Delta PE_{IBS}$ -0.00009 0.00008
scania 1 931 6 1 042 $\Delta PE_{IBS}$ -0.00001 -0.00005
TRACE 1 878 6 958 $\Delta PE_{IBS}$ 0.00001 0.00004
dataDIVAT2 1 837 4 508 $\Delta PE_{IBS}$ 0.00009 0.00001
actg 1 151 11 76 $\Delta PE_{IBS}$ -0.00000 0.00001
LeukSurv 1 043 7 441 $\Delta PE_{IBS}$ -0.00005 -0.00015
rdata 1 040 7 513 $\Delta PE_{IBS}$ -0.00013 0.00001
grace 1 000 5 110 $\Delta PE_{IBS}$ -0.00008 0.00032
Table 4: Difference in prediction error (Section 4.1) using benchmark datasets from SurvSet (Drysdale, 2022). For each dataset with at least 1 000 observations and no time-varying covariates, we fit a survival forest with identical options except for the splitting rule, and report prediction error. Some datasets have very low event rates, and the last column reports prediction error with deeper trees grown under less restrictive constraints on the number of events in each child node (grf option: alpha=0).